OpenStudy (anonymous):

Which of these values for P and a will cause the function to be an exponential growth function? A. P = ; a = 5 B. P = 4; a = C. P = 4; a = 1 D. P = ; a =

OpenStudy (anonymous):

sorry it didn't all show up, for(a) p=1/4 (b) a=1/5 for (d) p=1/4 a=1/5